### v1

@@ 1. Introduction
Let $G$ be a simply connected nilpotent Lie group and $\Gamma$ be a co-compact lattice so that $M=G/\Gamma$ is a nilmanifold equipped with the Haar measure $\mu$. In this paper, we study random walks by a finite set of translations on $M$. These random walks will be defined by an $m\in{\mathbb{N}},\ x\in M$, a set $F:=\{g_{1},\ldots,g_{m}\}\subset G$, and an associated probability vector $\vec{p}=(p_{1},p_{2},\dots,p_{m})$, i.e. $p_{i}>0$ and $\sum p_{i}=1$. The random walk is then a Markov chain where $x_{n}=g_{k}x_{n-1}$ with the probability $p_{k}$. Associated to the set $F$, we have an operator ${\mathcal{L}}:C^{r}(M)\rightarrow C^{r}(M)$ defined by
@@
It follows that for any $N>1$, $({\mathcal{L}}^{N}A)(x)={\mathbb{E}}_{x}(A(x_{N})).$ Since a random walk is exactly defined by its generators, we can view the product $\mu^{m}$ as a measure on the space of all random walks generated by $m$ translations. When we speak of the measure of a set of walks, we mean it with respect to this measure.
@@
As we shall see, the argument requires certain polynomials arising from the Lie algebra structure to be linearly independent.
@@
We now describe the structure of the paper. Section 2 contains preliminaries on nilpotent Lie groups, Lie algebras, and nilmanifolds. In Section 3, we define $m$-greatness and show that it implies rapid mixing for a full measure set of $m$-tuples modulo a lemma that connects $m$-greatness to Diophantine properties. In Section 4, we show the lemma by constructing words that act Diophantinely on appropriate tori. In Section 5, we establish that the groups listed in Theorem 1.3 are 22-great implying Theorem 1.3. We also show that any step $s$-nilmanifold is $s$-great, proving Theorem 1.2. Finally, we provide an example of a Lie algebra that is not $2$-great, showing that our technique is not sufficient to show the suspected optimal result that $N_{G}\!\!=\!\!2$ for all nilpotent groups.
@@ Acknowledgement
The authors greatly appreciate Adam Kanigowski for fruitful discussions and several suggestions for improving the quality of the initial draft. They also thank Bassam Fayad and Giovanni Forni for their several comments on the draft. M.K. thanks Xuesen Na for his assistance in setting up a computer program for testing calculations.
M.K. is grateful to the Institut de Mathématiques de Jussieu-Paris Rive Gauche and the Center of Excellence “Dynamics, Mathematical Analysis and Artificial Intelligence” at Nicolaus Copernicus University in Toruń for their hospitality during his visits.
@@ 2.1. Nilpotent Lie groups and Lie algebras
[4] If $G$ is a connected and simply connected nilpotent Lie group, then exponential map $\exp:\mathfrak{g}\rightarrow G$ is a diffeomorphism.
@@
If $\Gamma=\displaystyle\left\{\exp\left(\sum_{j,k}m_{j,k}X_{j}^{(k)}\right)\right\}_{m_{j,k}\in{\mathbb{Z}}}$, then we say that the basis is strongly based at $\Gamma$. The Lie algebra $\mathfrak{g}$ of any nilmanifold $G/\Gamma$ can be equipped with a Malcev basis strongly based at $\Gamma$. Moving forward, we will always use this basis when writing Lie algebra elements in coordinates. For convenience, we will denote by $X_{i}=X_{i}^{(0)}$.
@@ 2.2. Nilmanifolds and Fibration
is a nilmanifold. It follows that $\pi^{(k)}:M=G/\Gamma\rightarrow M^{(k)}=G/G^{(k+1)}\Gamma$ is a fibration whose fibers are the orbits of $G^{(k+1)}$ on $G/\Gamma$, homeomorphic to the nilmanifolds $G^{(k+1)}/(G^{(k+1)}\cap\Gamma)$. We can also define $M_{p}:=G^{(p)}/G^{(p+1)}\Gamma_{p}\simeq{\mathbb{T}}^{n_{p}}$.
@@ 2.3. Main examples
2.4. Harmonic Analysis on Nilmanifolds
In this section, we wish to describe the structure of $L^{2}(M)$ for a nilmanifold $M$. The first step is to describe the characters of $M$, denoted by $\widehat{M}$. For any Lie algebra $\mathfrak{a}$ let $\mathfrak{a}^{*}$ represent its dual as a vector space. Since we have a basis for $\mathfrak{g}$ (the Malcev basis), the coordinate functionals $X_{1}^{(1)*},\ldots,X_{n_{s-1}}^{(s-1)*}$ are well-defined and form a basis for $\mathfrak{g}^{*}$. Define $\pi_{p}:\mathfrak{g}\rightarrow{\mathbb{R}}^{n_{p}}$ by $\pi_{p}(V)=(X_{1}^{(p)*}(V),\ldots,X_{n_{p}}^{(p)*}(V))$. Also, let $E_{p}^{*}=\{X_{1}^{(p)*},\ldots,X_{n_{p}}^{(p)*}\}$, and define $\Lambda_{p}$ to be the set of integer linear combinations of elements of $E_{p}^{*}$. We let $\Lambda=\bigoplus_{i=0}^{s-1}\Lambda_{i}.$ Precomposing an element of $\Lambda$ with $\log$ and postcomposing by the complex exponential $e(x)=e^{2\pi ix}$ gives a character on $M$. In fact, $\widehat{M}$ is exactly equal to $e\circ\Lambda\circ\log$. When $\lambda\in\Lambda$, we will let $\chi_{\lambda}$ denote the character defined by $\chi_{\lambda}(g)=e^{2\pi i\lambda(\log g)}$. We define a a norm on $\widehat{M}$ by $\|\chi_{\lambda}\|=\|\lambda^{*}\|$ where $\lambda^{*}$ is the element of $\mathfrak{g}$ dual to $\lambda$. Finally, for any characters $\chi\in\widehat{M}$, we let
$H_{\chi}=\{\varphi\in L^{2}(M):\varphi(gx)=\chi(g)\varphi(x)\}.$
We have that
$L^{2}(M)=\bigoplus_{\chi\in\widehat{M}}H_{\chi}.$
@@ 2.5. Diophantine Conditions
This result will provide the bridge from the Lie algebra structure to the Diophantine estimates needed to show mixing. Since the functions we will consider will always be homogeneous polynomials with positive degree, linear independence of $f_{1},...,f_{n}$ is the only condition we will need to check in order to apply this theorem in practice.
@@ 3. $m$-greatness and mixing
In this section, we will first define what it means for a nilmanifold to be $m$-great and then show that on these manifolds almost any $m$-tuple generates a rapidly mixing walk.
@@ 3.1. Linear independence of polynomials
$\mathcal{D}=\{\bar{t}\in{\mathbb{R}}^{a}:\mathcal{H}_{1}(\bar{t},\alpha),\dots\mathcal{H}_{\ell}(\bar{t},\alpha)\text{ are linearly dependent}\}.$
Expanding $\mathcal{H}_{j}(\bar{t},\alpha)=\sum_{{\mathbf{m}}}c_{j,{\mathbf{m}}}(\bar{t})\alpha^{{\mathbf{m}}}$ we see that the above conditions amounts to vanishing of certain minors of the matrix $c_{j,{\mathbf{m}}}$ whence $\mathcal{D}$ is an algebraic subvariety of ${\mathbb{R}}^{a}.$ We say that $\mathcal{H}$ is non-degenerate if $\mathcal{D}\neq{\mathbb{R}}^{a}.$
Now we describe special polynomial mapping associated to a Lie algebra. Fix integers $m$ and $p.$ Consider $m$ vectors $V_{i}=\sum_{j=1}^{n}\alpha_{ij}X_{j},$ $i\in\{1,\dots m\}$ and let
$\mathcal{H}_{m,p}(t,\alpha)=[\ldots\bf{[t_{0}V,t_{1}V],\ldots],t_{p}V]}+\mathfrak{g}^{(p+1)}\in\mathfrak{g}^{(p)}/\mathfrak{g}^{(p+1)}$
where ${\mathbf{t}}_{q}=(t_{q1},\dots t_{qm})$ and ${\mathbf{t}}\textbf{V}=\sum_{i=1}^{m}t_{i}V_{i}.$ Note that we have
@@
$\mathcal{H}_{m,p}({\mathbf{k}},\alpha)=\sum_{i=1}^{n_{p}}P_{i}(k,\alpha)X_{i}^{(p)}+\mathfrak{g}^{(p+1)}$
where each $P_{i}({\mathbf{k}},\alpha)$ is a polynomial in the coordinates of $\mathbf{k}$} and $\alpha_{ij}$. From now on we will suppress the $\mathfrak{g}^{(p+1)}$ that ought to appear whenever we discuss $\mathcal{H}_{2,p}$, but we agree that $\mathcal{H}_{2,p}$ is always defined modulo $\mathfrak{g}^{(p+1)}$. We now provide a brief illustrative example.
@@
We clearly see that $\mathcal{H}_{2,2}$ is non degenerate since even just setting $k_{01}=k_{12}=k_{21}=1$ and $k_{02}=k_{11}=k_{22}=0$, we get that at that value of $\mathbf{k}$} $\mathcal{H}_{2,2}$ evaluates to $M_{121}$, which has linearly independent rows.
@@ 3.2. $m$-great Lie algebras and mixing
If for every $i\in[1,s-1]$, the $\mathcal{H}_{m,i}$ is non degenerate we say that $\mathfrak{g}$ is mm–great.
@@
Let $M=G/\Gamma$ be a nilmanifold and $(g_{1},\ldots g_{m})$ be an $m$-tuple.
  • A pair of words $W_{1},W_{2}\in\langle g_{1},\ldots,g_{m}\rangle$ is said to be nice on $G^{(p)}$ if they have equal length and their difference $h:=W_{2}W_{1}^{-1}$ satisfies $h\in G^{(p)}$, and there exist $\gamma,\tau>0$ such that $\pi_{p}(\log(h))\in DC(\gamma,\tau)$.
@@
To prove this Theorem 3.5, we will show that for any nontrivial character $\chi$, $\mathcal{L}$} has spectral gap on $H_{\chi}$, and, moreover, that the size of the spectral gap has a polynomial lower bound as $\chi$ grows in norm. We will also use the fact that the size of a $C^{r}$ function’s projection into $H_{\chi}$ is polynomially bounded in the norm of $\chi$. Together, these estimates will allow us to prove Theorem 3.5. We now show the spectral gap.
@@
Assume that $(g_{1},...,g_{m})$ acts Diophantinely at all levels. There exist constants $C_{1},C_{2},\ell,\tau>0$ such that if $\chi$ is a nontrivial character on $M$ and $A\in H_{\chi}$, then
@@
Let $\lambda=\lambda_{0}\oplus\ldots\oplus\lambda_{s-1}\in\Lambda$ be such that $\chi=\chi_{\lambda}$. Let $i$ be the largest value such that $\lambda_{i}\neq 0$. Let $W_{1}$ and $W_{2}$ be nice on $G^{(i)}$. Let $\ell_{i}$ be the common length of $W_{1}$ and $W_{2}$, and let $\gamma_{i},\tau_{i}$ be the associated Diophantine constants. We note that each of these constants depend only on $i$ and not on the particular character $\chi$. Let ${\mathbf{p}}=\min(p(W_{1}),p(W_{2})).$ Suppose without loss of generality that ${\mathbf{p}}=p(W_{1})$. Then setting $h=W_{2}W_{1}^{-1}$
@@
By Lemma A.1 from the appendix, we have that the first term satisfies
${\mathbf{p}}|A(W_{1}x)+A(hW_{1}x)|={\mathbf{p}}|1+\chi(h)||A(W_{1}x)|\leq 2{\mathbf{p}}(1-{\bar{C}}_{i}\|\chi\|^{-2\tau_{i}})\|A\|_{C^{0}},$
while the second term is at most $(1-2{\mathbf{p}})\|A\|_{C^{0}}.$ Combining these estimates, we see that on $H_{\chi}$,
$\|{\mathcal{L}}^{\ell_{i}}A\|_{C^{0}}\!\!\leq\!\!(1\!-\!{\hat{C}}_{i}\|\chi\|^{-2\tau_{i}})\|A\|_{C^{0}}.$
Fixing ${\hat{C}}=\min\{{\hat{C}}_{i}\}$, $\ell=\max\{\ell_{i}\}$, and $\tau=\max\{\tau_{i}\}$, we see that
@@
$\sum_{\|\chi\|^{2\tau}\geq\sqrt{N}}\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}\leq\|A\|_{C^{r}}\sum_{\|\chi\|>N^{1/4\tau}}\frac{1}{\|\chi\|^{r}}\leq C_{4}N^{-(r-\dim(M))/4\tau}\|A\|_{C^{r}}.$
@@
Whenever an $m$-tuple acts Diophantinely on every level, Theorem 3.5 applies. By (3.3), this is enough to conclude rapid mixing for the random walk they generate. Lemma 3.4, implies that the set of such tuples has full measure with respect to $\mu^{m}$. ∎

### v2

@@ 1. Introduction
Let $G$ be a simply connected nilpotent Lie group and $\Gamma$ be a co-compact lattice so that $M=G/\Gamma$ is a nilmanifold equipped with the Haar measure $\mu$. A translation on $M$ by $g\in G$ is the map $x\Gamma\mapsto gx\Gamma$. In this paper, we study random walks by a finite set of translations on $M$. These random walks will be defined by an $m\in{\mathbb{N}},\ x\in M$, a set $F:=\{g_{1},\ldots,g_{m}\}\subset G$, and an associated probability vector $\vec{p}=(p_{1},p_{2},\dots,p_{m})$, i.e. $p_{i}>0$ and $\sum p_{i}=1$. The random walk is then a Markov chain where $x_{n}=g_{k}x_{n-1}$ with the probability $p_{k}$. We will let $\langle F\rangle$ denote the semigroup generated by the set $F$. Associated to the set $F$, we have an operator ${\mathcal{L}}:C^{r}(M)\rightarrow C^{r}(M)$ defined by
@@
It follows that for any $N>1$,
$({\mathcal{L}}^{N}A)(x)={\mathbb{E}}_{x}(A(x_{N}))=\sum_{|W|=N}p_{N}(W)A(Wx),$
where $|W|$ is the length of the word $W$ and ${p_{N}(W)=p_{N}(g_{w_{N}}\ldots g_{w_{1}})}=\prod_{i=1}^{N}p_{w_{i}}$ is the probability of having walked by $W$ at time $N$. Since a random walk is exactly defined by its generators, we can view the product $\mu^{m}$ as a measure on the space of all random walks generated by $m$ translations. When we speak of the measure of a set of walks, we mean it with respect to this measure.
@@
It is a classical theorem that for single translations on a nilmanifold, ergodicity, unique ergodicity, and minimality are equivalent and in fact only depend on the irrationality of the abelian projection of the translation. Our result is similar in that we only need an arithmetic property on the restriction of the action in the maximal torus factor. On the other hand, unlike other cases, the arithmetic property we require takes into account how higher-order commutators of the generators will act on $M$. In particular, the argument requires certain polynomials arising from the Lie algebra structure to be linearly independent. In fact, the relevant polynomials only depend on the coordinates of the generators in the abelianization of $G$, and the full measure set of rapid mixing walks we find corresponds to a full measure set of abelian parts that act Diophantinely in some sense (see Definition 3.3). It follows that Theorem A could be strengthened to say that any $m$-tuples in $G$ that project into the aforementioned full measure set of abelian parts generate rapid mixing walks.
@@
We now describe the structure of the paper. Section 2 contains preliminaries on nilpotent Lie groups, Lie algebras, and nilmanifolds. In Section 3, we define $m$-greatness as well as a relevant Diophantine property. We show that the Diophantine property implies rapid mixing (Proposition 3.7). Then, modulo a claim that connects $m$-greatness to the Diophantine property (Proposition 3.4), we show that a full measure set of $m$-tuples on an $m$-great nilmanifold generate rapidly mixing walks. In Section 4, we show Proposition 3.4 by constructing words that act Diophantinely on appropriate tori. In Section 5, we establish that the groups listed in Theorem B are 22-great implying Theorem B. We also show Proposition 5.5 which states that any step $s$-nilmanifold is $s$-great, proving Theorem A. Finally, we provide an example of a Lie algebra that is not $2$-great, showing that our technique is not sufficient to show the suspected optimal result that $N_{G}\!\!=\!\!2$ for all nilpotent groups.
@@ Acknowledgement
The authors appreciate Adam Kanigowski for fruitful discussions and several suggestions for improving the quality of the initial draft. They are grateful to Bassam Fayad and Giovanni Forni for their careful readings and several comments on the draft. M.K. thanks Xuesen Na for his assistance in setting up a computer program for testing calculations.
@@ 2.1. Nilpotent Lie groups and Lie algebras
[4] If $G$ is a connected and simply connected nilpotent Lie group, then the exponential map $\exp:\mathfrak{g}\rightarrow G$ is a diffeomorphism.
@@
If $\Gamma=\displaystyle\left\{\exp\left(\sum_{j,k}m_{j,k}X_{j}^{(k)}\right)\right\}_{m_{j,k}\in{\mathbb{Z}}}$, then we say that the basis is strongly based at $\Gamma$. The Lie algebra $\mathfrak{g}$ of any nilmanifold $G/\Gamma$ can be equipped with a Malcev basis strongly based at $\Gamma$. Moving forward, we will always use this basis when writing Lie algebra elements in coordinates. For convenience, we will denote by $X_{i}=X_{i}^{(0)}$. We observe that if $A\subseteq{\mathbb{R}}^{n_{0}}$, then
$\mu(\{\alpha_{01}X_{1}+\ldots\alpha_{0n_{0}}X_{n_{0}}+O(\mathfrak{g}_{1}):(\alpha_{01},\ldots,\alpha_{0n_{0}})\in A\}=\mu(A),$
i.e. for a set defined by the coefficients of the $X_{i}$, the measure of the set is exactly the Lebesgue measure of the permitted coefficients.
@@ 2.2. Nilmanifolds and Fibration
is a nilmanifold. It follows that $\pi^{(k)}:M=G/\Gamma\rightarrow M^{(k)}$ is a fibration whose fibers are the orbits of $G^{(k+1)}$ on $G/\Gamma$, homeomorphic to the nilmanifolds $G^{(k+1)}/(G^{(k+1)}\cap\Gamma)$. We can also define $M_{p}:=G^{(p)}/G^{(p+1)}\Gamma_{p}\simeq{\mathbb{T}}^{n_{p}}$.
2.3. Harmonic Analysis on Nilmanifolds
In this section, we wish to describe the structure of $L^{2}(M)$ for a nilmanifold $M$. To do this, we will need to describe the characters of $M$, denoted by $\widehat{M}$. For any Lie algebra $\mathfrak{a}$, let $\mathfrak{a}^{*}$ represent its dual as a vector space. Since we have a basis for $\mathfrak{g}$ (the Malcev basis), the coordinate functionals $X_{1}^{(1)*},\ldots,X_{n_{s-1}}^{(s-1)*}$ are well-defined and form a basis for $\mathfrak{g}^{*}$. Define $\pi_{p}:\mathfrak{g}\rightarrow{\mathbb{R}}^{n_{p}}$ by $\pi_{p}(V)=(X_{1}^{(p)*}(V),\ldots,X_{n_{p}}^{(p)*}(V))$. Also, let $E_{p}^{*}=\{X_{1}^{(p)*},\ldots,X_{n_{p}}^{(p)*}\}$, and define $\Lambda_{p}$ to be the set of integer linear combinations of elements of $E_{p}^{*}$.
[…]
$H_{\chi}=\{\varphi\in L^{2}(M):\varphi(gx)=\chi(g)\varphi(x)\}.$
We have that
$L^{2}(M)=\bigoplus_{\chi\in\widehat{M}}H_{\chi}.$
@@ 2.5. Diophantine Conditions
This result will provide the bridge from the Lie algebra structure to the Diophantine estimates needed to show mixing. Since the functions we will consider will always be homogeneous polynomials with positive degree, linear independence of $f_{1},\ldots,f_{n}$ is the only condition we will need to check in order to apply this theorem in practice.
@@ 3. $m$-greatness and mixing
In this section, we will first define what it means for a nilmanifold to be $m$-great and then show that on these manifolds, almost any $m$-tuple generates a rapidly mixing walk.
@@ 3.1. Linear independence of polynomials
$\mathcal{D}(\mathcal{H})=\{\bar{t}\in{\mathbb{R}}^{a}:\mathcal{H}_{1}(\bar{t},\alpha),\dots,\mathcal{H}_{\ell}(\bar{t},\alpha)\text{ are linearly dependent}\}.$
Expanding $\mathcal{H}_{j}(\bar{t},\alpha)=\sum_{{\mathbf{m}}}c_{j,{\mathbf{m}}}(\bar{t})\alpha^{{\mathbf{m}}}$ we see that the above conditions amounts to vanishing of certain minors of the matrix $c_{j,{\mathbf{m}}}$ whence $\mathcal{D}(\mathcal{H})$ is an algebraic subvariety of ${\mathbb{R}}^{a}.$ We say that $\mathcal{H}$ is non-degenerate if $\mathcal{D}(\mathcal{H})\neq{\mathbb{R}}^{a}.$
Now we describe a special polynomial mapping associated to a Lie algebra. Fix integers $m$ and $p.$ Consider $m$ vectors $V_{i}=\sum_{j=1}^{n}\alpha_{ij}X_{j}$ with $i\in\{1,\ldots,m\}$. These will correspond to the exponential coordinates of $m$ generators of a random walk $g_{1},\ldots,g_{m}$ mod $\mathfrak{g}^{(1)}$. Thus, each variable $\alpha_{ij}$ corresponds to the $X_{j}$ coordinate of $g_{i}$. The idea is to see in coordinates what group elements will arise as we take iterated brackets. To that effect, let
$\mathcal{H}_{m,p}({\mathbf{k}},\alpha)=[\ldots\bf{[k_{0}V,k_{1}V],\ldots],k_{p}V]}+\mathfrak{g}^{(p+1)}\in\mathfrak{g}^{(p)}/\mathfrak{g}^{(p+1)},$
where ${\mathbf{k}}=({\mathbf{k}}_{0},\ldots,{\mathbf{k}}_{p})$, ${\mathbf{k}}_{q}=(k_{q1},\dots,k_{qm})$ and ${\mathbf{k}}_{q}\textbf{V}=\sum_{i=1}^{m}k_{qi}V_{i}.$
The image of $\mathcal{H}_{m,p}$ is exactly the $\mathfrak{g}^{(p)}$ coordinates of the elements $p$-fold brackets of vectors in the span of the $V_{i}$. As we shall see, this will bear relation to the elements $G$ achievable as $p$-fold brackets of elements of the group generated by the $g_{i}$. We use $\mathbf{k}$} instead of $t$ for the first parameter of this map because, while it still makes sense to evaluate this map at any real time, we will be particularly interested in $\mathcal{H}_{m,p}({\mathbf{k}},\alpha)$ that occurs as differences between two different words in the walk. These differences correspond to special values of $\mathbf{k}$}, where $\mathbf{k}$} counts how many times each generator was used at each step of the construction of the words used to obtain the difference $\mathcal{H}_{m,p}({\mathbf{k}},\alpha)$, and thus will always be integral. Note that we have
@@
$\mathcal{H}_{m,p}({\mathbf{k}},\alpha)=\sum_{\ell=1}^{n_{p}}P_{\ell}({\mathbf{k}},\alpha)X_{i}^{(p)}+\mathfrak{g}^{(p+1)}$
where each $P_{\ell}({\mathbf{k}},\alpha)$ is a polynomial in the coordinates of $\mathbf{k}$} and $\alpha_{ij}$ (in particular the $\alpha_{0j}$). We will care about when the map $\mathcal{H}_{m,p}$ is degenerate, so in context we have that $\mathcal{D}(\mathcal{H}_{m,p})$ represents the variety
$\{{\bf k}\in{\mathbb{R}}^{m(p+1)}:P_{1}({\bf k},\alpha),\dots,P_{n_{p}}({\bf k},\alpha)\text{ are linearly dependent polynomials in }\alpha\}.$
From now on we will suppress the $\mathfrak{g}^{(p+1)}$ that ought to appear whenever we discuss $\mathcal{H}_{2,p}$, but we agree that $\mathcal{H}_{2,p}$ is always defined modulo $\mathfrak{g}^{(p+1)}$. We now provide a brief illustrative example.
@@
We clearly see that $\mathcal{H}_{2,2}$ is non-degenerate since even just setting $k_{01}=k_{12}=k_{21}=1$ and $k_{02}=k_{11}=k_{22}=0$, we get that at that value of $\mathbf{k}$}, $\mathcal{H}_{2,2}$ evaluates to $M_{121}$, which has linearly independent rows. Thus, our selected value of $\bf k$} is not in $\mathcal{D}(\mathcal{H}_{2,2})$, and thus $\mathcal{H}_{2,2}$ is non-degenerate.
@@ 3.2. $m$-great Lie algebras and mixing
A Lie algebra $\mathfrak{g}$ is $m$-great if $\mathcal{H}_{m,i}$ is non-degenerate for every $i\in[0,s-1]$.
@@
Let $M=G/\Gamma$ be a step-$s$ nilmanifold and $F=(g_{1},\ldots,g_{m})$ be an $m$-tuple.
  • A pair of words $W_{1},W_{2}\in\langle F\rangle$ is said to be nice on $G^{(p)}$ if they have equal length and their difference $h:=W_{2}W_{1}^{-1}$ satisfies $h\in G^{(p)}$, and there exist $\gamma,\tau>0$ such that $\pi_{p}(\log(h))\in DC(\gamma,\tau)$.
@@
To prove Proposition 3.5, we will show that for any nontrivial character $\chi$, $\mathcal{L}$} has spectral gap on $H_{\chi}$, and, moreover, that the size of the spectral gap has a polynomial lower bound in the norm of $\chi$. We will also use the fact that the size of a $C^{r}$ function’s projection into $H_{\chi}$ is polynomially bounded in the norm of $\chi$. Together, these estimates will allow us to prove Theorem 3.5. We now show the spectral gap.
@@
Assume that $F$ acts Diophantinely at all levels. There exist constants $C_{1},C_{2},\ell,\tau>0$ such that if $\chi$ is a nontrivial character on $M$ and $A\in H_{\chi}$, then
@@
Let $A\in H_{\chi}$ where $\chi\in\widehat{M}^{(i)}$. We will now fix words and constants that depend only on $i$ and not on the particular character $\chi$ as follows. Let $W^{i}_{1}$ and $W^{i}_{2}$ be the nice words on $G^{(i)}$ guaranteed by the assumption that the $g_{i}$ act Diophantinely at all levels. Let $\ell_{i}$ be the common length of $W^{i}_{1}$ and $W^{i}_{2}$, and let $\gamma_{i},\tau_{i}$ be the associated Diophantine constants.
Recall that $p_{\ell_{i}}(W)$ denotes the probability that the walk has moved by $W$ at time $\ell_{i}$. Suppressing the dependence on $\ell_{i}$, let ${\mathbf{p}}=\min(p(W_{1}),p(W_{2}))$. Suppose without loss of generality that ${\mathbf{p}}=p(W_{1})$. Then set $h=W_{2}W_{1}^{-1}$, and let $W_{3},\ldots,W_{m^{\ell_{i}}}$ denote the remaining words with positive probability at step $\ell_{i}$ of the walk. With this notation, we have
@@
Since $A\in H_{\chi}$, we have that
${\mathbf{p}}\left[A(W_{1}x)+A(hW_{1}x)\right]={\mathbf{p}}\left[A(W_{1}x)+\chi(h)A(W_{1}x)\right]=(1+\chi(h))A(W_{1}x).$
Letting $\lambda\in\Lambda_{p}$ to be the nonzero functional such that $\chi=\chi_{\lambda}$, we get that
$|1+\chi(h)|=|1+e^{2\pi i\lambda(\log(h))}|.$
Since by niceness of $W_{1}^{i}$ and $W_{2}^{i}$, $\log(h)\in DC(\gamma,\tau)$, we can apply Lemma A.2 from the appendix, to say that, for some constant $\bar{C}_{i}$ depending only on $\gamma_{i}$, we have that
$|1+\chi(h)|\leq 2-\bar{C}_{i}|\chi|^{-2\tau_{i}}.$
Putting this all together, we find that
${\mathbf{p}}|A(W_{1}x)+A(hW_{1}x)|={\mathbf{p}}|1+\chi(h)||A(W_{1}x)|\leq 2{\mathbf{p}}(1-{\bar{C}}_{i}\|\chi\|^{-2\tau_{i}})\|A\|_{C^{0}}.$
Turning our attention to the second term in the expression for ${\mathcal{L}}^{\ell_{i}}(A)(x),$ we see that
$(p(W_{2})-p(W_{1}))A(W_{2}x)+\sum_{j\geq 3}p(W_{j})A(W_{j}x)\leq(1-2{\mathbf{p}})\|A\|_{C^{0}}.$
Combining these estimates, we see that on $H_{\chi}$,
$\|{\mathcal{L}}^{\ell_{i}}A\|_{C^{0}}\leq 2{\mathbf{p}}(1-{\bar{C}}_{i}\|\chi\|^{-2\tau_{i}})\|A\|_{C^{0}}+(1-2{\mathbf{p}})\|A\|_{C^{0}}\leq(1-{\hat{C}}_{i}\|\chi\|^{-2\tau_{i}})\|A\|_{C^{0}},$
where $\hat{C}_{i}=2{\mathbf{p}}\bar{C}_{i}$. We have now shown the desired result except for an unwanted dependence on the level $i$ in the constants $\ell_{i},\hat{C}_{i}$, and $\tau_{i}$. Since there are only $s$ different levels, we only used $s$ different word pairs with associated constants. Thus, we can simply fix ${\hat{C}}=\min\{{\hat{C}}_{i}\}$, $\ell=\max\{\ell_{i}\}$, and $\tau=\max\{\tau_{i}\}$. We then see that
@@
$\sum_{\|\chi\|^{2\tau}\geq\sqrt{N}}\|{\mathcal{L}}^{N}A_{\chi}\|_{C^{0}}\leq\sum_{\|\chi\|>N^{1/4\tau}}\frac{\|A\|_{C^{r}}}{\|\chi\|^{r}}\leq C_{4}N^{-(r-\dim(M))/4\tau}\|A\|_{C^{r}}.$
@@
Whenever an $m$-tuple acts Diophantinely on every level, Proposition 3.5 applies. By (3.3), this is enough to conclude rapid mixing for the random walk they generate. Proposition 3.4 implies that the set of such tuples has full measure with respect to $\mu^{m}$. ∎